if 2x+y=7 and 3x-2y=3,
by how much is 7x greater than 10?

Answer by Cromlix(4381)

2x + y = 7......1
3x - 2y = 3.....2
Multiply (1) by 2
4x + 2y = 14.....3
3x - 2y = 3......2
Add (3) + (2)
7x = 17
x = 17/7
Substitute x = 17/7 into:
2x + y = 7
34/7 + y = 49/7 (7)
y = 49/7 - 34/7
y = 15/7.
7x = 7 * 17/7 = 17
7x is 7 units greater than 10.
Hope this helps.
:-)